OpenStudy (anonymous):

Find the limit of xln(x) as x approaches 0 from the right. (Use L'Hopital's Rule)

OpenStudy (anonymous):

It can be re-written as ln(x)/(1/x), but isn't that -inf/inf? Can you use L'H's rule when the signs of infinity are different?

OpenStudy (anonymous):

Yes. That's one of the indeterminate forms.

OpenStudy (anonymous):

\[\frac{0}{0},\pm\frac{\infty}{\infty}\]

OpenStudy (anonymous):

Oh, okay, easy enough.